OpenStudy (anonymous):

solve using (V.E.S.T-Variable, Equation, Solve, Try it.) In a triangle, the middle lenght side is 3 more the short side the largest is double the middle side the perimeter is 45cm What are the lengths of the sides

OpenStudy (anonymous):

Start by setting some variables. Let's call the shortest side a, the middle side b, and the longest side c. How can you rephrase the first two conditions (the middle side is 3 more than the short side, and the largest side is double the middle side) in terms of those three variables?

OpenStudy (anonymous):

a+3=b b(2)=c

OpenStudy (anonymous):

Right. Now what is the third condition in terms of the three variables? The perimeter being 45cm.

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

That was a question haha, how would you write the perimeter being 45 cm in terms of the three variables?

OpenStudy (anonymous):

thats mostly were i was confused

OpenStudy (anonymous):

What is the perimeter of a triangle? What does that mean?

OpenStudy (anonymous):

The outer limits of an area.

OpenStudy (anonymous):

The perimeter is the sum of all of the sides of the figure. Using that definition, what's the formula in terms of the variables that we have?

OpenStudy (anonymous):

a+b+c=45cm

OpenStudy (anonymous):

Exactly. Now, the next part is the only part that's a little tricky. We want to phrase everything in terms of one variable. We have b=a+3 already, and c=2b. From those we can get that c=2a+6 by substituting b=a+3 into the second formula. So, we can rewrite a+b+c as a+(a+3)+(2a+6). So, a+(a+3)+(2a+6)=45. Can you solve it from there?

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

um no im not getting it yet like how do i get rid of the a b an c

OpenStudy (anonymous):

Well, I already got rid of the b and c for you in the last comment. If a+(a+3)+(2a+6)=45, what is a?

OpenStudy (anonymous):

45 divided by 3 =15 a=9 b=12 c=24